# IPPP/24/10

### v1

@@ Abstract
We delve into the use of photonic quantum computing to simulate quantum mechanics and extend its application towards quantum field theory. We develop and prove a method that leverages this form of Continuous-Variable Quantum Computing (CVQC) to reproduce the time evolution of quantum-mechanical states under arbitrary Hamiltonians, and we demonstrate the method’s remarkable efficacy with various potentials. Our method centres on constructing an evolver-state, a specially prepared quantum state that induces the desired time-evolution on the target state. This is achieved by introducing a non-Gaussian operation using a measurement-based quantum computing approach, enhanced by machine learning. Furthermore, we propose a novel framework in which these methods can be extended to encode field theories in CVQC without discretising the field values, thus preserving the continuous nature of the fields. This opens new avenues for quantum computing applications in quantum field theory.
@@ 1 Introduction
Harnessing the intricate dynamics of quantum mechanics to improve our understanding of fundamental physics has lead to the pursuit of computational paradigms that go beyond classical boundaries and allow for complex systems to be simulated with devices which themselves are inherently quantum mechanical. Such devices would leverage the properties of quantum systems, such as superposition and entanglement, to perform calculations directly using the intrinsic dynamics of the system. Two main paradigms have been identified: digital and analogue quantum computing. The former encodes information onto systems with a finite number of discrete degrees of freedom, such as the qubit, a two-state quantum system. The latter instead encodes information on systems which are described by operators which have a continuous spectrum. This paradigm is often called Continuous-Variable Quantum Computing (CVQC) Lloyd1999 ; RevModPhys.77.513 ; Ladd2010 ; Adesso_2014 . Among the various kinds of CVQC, quantum optics emerges as a fascinating framework, using the infinite-dimensional landscape of photon states to encode and manipulate information PhysRevA.73.032318 ; Kok2007 ; Slussarenko2019 ; Killoran_2019 ; Bromley_2020 ; Eaton_2022 ; Taballione2023modeuniversal . The eigenstates of these operators form an infinite-dimensional Hilbert space, with continuous-variable analog of the qubit being the so-called qumode.
@@
CVQC, rooted in the continuous spectra of quantum operators, offers various possibilities for simulating the dynamics of quantum particles and fields. By employing such qumodes — quantum analogs of classical harmonic oscillators — as the fundamental information units, CVQC provides a natural framework for encoding quantum states in the continuous observables of photons, e.g. their position or momentum. This paradigm allows for implementing Gaussian gate operations, which manipulate the quantum states through transformations that preserve their Gaussian character, thereby enabling a broad range of quantum simulations. However, the true power of CVQC unfolds with the inclusion of non-Gaussian operations, which introduce higher-order interactions essential for achieving universal quantum computation Lloyd1999 ; PhysRevA.100.012326 ; PhysRevA.100.052301 . These operations, albeit challenging to implement due to the weakly interacting nature of photons, open the door to simulating complex quantum systems with high fidelity. Most existing proposals to simulate quantum systems on CVQC rely on specific ways to induce non-Gaussian effects, such as the Kerr effect. However, the non-Gaussian characteristics introduced by current non-linear optical materials are very weak RevModPhys.77.513 ; Slussarenko2019 ; Kok2007 ; Ladd2010 ; PhysRevA.100.052301 , and constructing an arbitrary non-Gaussian operation is difficult. An alternative approach is achieved by integrating measurement-based quantum computing techniques PhysRevLett.86.5188 ; Booth:2021hvw and leveraging the entanglement of qumodes. Through this, it is possible to instead induce the desired non-Gaussian characteristics, paving the way for simulations that capture non-trivial quantum dynamics.
@@
A central aspect of quantum systems that one might wish to explore using such methods is the Hamiltonian and the time-evolution that is governed by it, a cornerstone in understanding the dynamics of quantum particles and fields Jordan:2011ci ; Jordan:2012xnu ; Jordan:2014tma ; PhysRevA.92.063825 ; Jordan_2018 ; Klco:2018zqz ; Banuls:2019bmf . By simulating the time evolution governed by a system’s Hamiltonian, we can explore how quantum states change over time, which is crucial for predicting the behaviour of quantum particles and systems under various conditions. This process is essential for simulating inherently quantum mechanical phenomena that cannot be accurately modelled using classical physics. One of the most notable examples is quantum tunnelling, a phenomenon where particles pass through potential barriers that would be insurmountable, according to classical mechanics. This process is critical in a wide range of quantum systems, from the decay of atomic nuclei to the operation of quantum dots and superconducting qubits. By simulating the time evolution of quantum systems, one can also investigate other observables, such as energy spectra, correlation functions, and phase transitions, providing deep insights into the nature of quantum materials, chemical reactions, and even the evolution of early universe conditions in high-energy physics. CVQC offers a novel approach to simulating the time evolution of quantum states under arbitrary Hamiltonians. By decomposing the time evolution into discrete steps through Trotterization, we show how CVQC facilitates the simulation of complex quantum systems, including those governed by non-Gaussian potentials.
@@
In contrast, quantum gate computing operates within a digital framework, encoding quantum information in discrete qubits and manipulating it through a sequence of quantum gates. While this approach has paved the way for significant advancements in quantum computing, and has gained interest for uses in the simulation of quantum field theories Jordan:2011ci ; Jordan:2012xnu ; Jordan:2014tma ; PhysRevA.92.063825 ; Jordan_2018 ; Klco:2018zqz ; Banuls:2019bmf ; Abel:2020qzm ; Abel:2020ebj ; PRXQuantum.4.027001 ; Kane:2022ejm , high-energy particle collisions Blance:2020ktp ; Bepari:2020xqi ; Bepari:2021kwv ; Gustafson:2022dsq , and machine learning Schuld2015 ; Carleo:2019ptp ; Abel:2022lqr ; Abel:2023erv ; Rousselot:2023pcj , it inherently approximates the continuous nature of quantum systems, potentially limiting its ability to capture the full spectrum of quantum dynamics. Thus, CVQC, with its continuous variable approach, offers a promising avenue for simulating quantum systems complementary to quantum gate computing approaches.
@@
The simulation of non-trivial Hamiltonians on CVQC devices has been difficult due to the challenges in implementing non-Gaussian operations on photonic devices. Most current approaches propose circuits which involve non-Gaussian gate operations generated by non-linear optics to achieve Hamiltonian simulation on a continuous-variable device PhysRevA.92.063825 ; PhysRevA.97.062311 . However, experimentally such operations are difficult to produce Lloyd1999 ; RevModPhys.77.513 ; Adesso_2014 . This has lead to attempts to generate non-Gaussian effects through the use of Gaussian operations and measurements PhysRevA.105.012412 , in particular References PhysRevA.100.012326 ; PhysRevA.100.052301 utilise a machine-learning routine to enhance the production of non-Gaussian states. In this paper, we build and improve on this approach and propose a novel quantum circuit for simulating the Trotterised time-evolution of a quantum-mechanical state under the influence of an arbitrary Hamiltonian using only Gaussian operations and measurements. The circuit for a single Trotter step can then be applied iteratively to achieve the time-evolution of the Hamiltonian. By using a top-hat resource function and generating the evolver-state, the non-Gaussian part of the Trotter evolution, using a measurement-based circuit, we show that the affect of the noise-factor from Reference PhysRevA.100.012326 can be maximally suppressed without reducing the strength of the overall operation. We demonstrate the circuit’s ability to simulate time-evolution for several examples of quantum-mechanical Hamiltonians and find that the circuit performs remarkably well compared to exact, classical simulations. Furthermore, we show that the approach can be extended, outlining how the continuous behaviour of the CVQC device can be harnessed to accurately simulate quantum field theories, maintaining the continuous nature of fields.
@@ 2 Background: Continuous-Variable Quantum Computing
The framework we will consider is quantum computation via quantum optics, which offers an experimentally realisable approach to Continuous-Variable Quantum Computing (CVQC), in which the continuous-variable system is constructed from the quantised electromagnetic field RevModPhys.77.513 ; Adesso_2014 . The use of quantum optics as a method of quantum computing benefits from the exceedingly low decoherence properties of photons, and therefore holds good potential for transmitting and maintaining quantum information throughout many gate operations Kok2007 ; Ladd2010 ; Bromley_2020 . In this realisation, the Hilbert space of each qumode is the infinite-dimensional photon-number degree of freedom EISERT2003 which is essentially the Fock basis, such that the total Hilbert space, $\mathscr{H}$, of an $N$ qumode system is
@@
$[\hat{x}_{i},\hat{p}_{j}]=i\hbar\delta_{ij}$ (4)
For the rest of this paper, we will adopt natural units, $\hbar=m=\omega=1$.
@@
where $i,j$ label the qumodes on which they act, and $\theta_{ij},\theta^{\prime}_{ij},\theta^{\prime\prime}_{ij}$ are constants. For universal computation it is essential also to be able to implement non-Gaussian gates (i.e. with $\hat{x}^{3}$ and higher appearing in the phase) Lloyd1999 ; PhysRevA.100.012326 ; PhysRevA.100.052301 . This is a delicate process because the fact that photons do not strongly interact with each other, while being excellent for maintaining coherence, is accompanied by the negative implication that non-Gaussian effects are hard to achieve. Directly producing non-Gaussian gate operations requires the use of non-linear optical materials which induce non-Gaussian effects, such as the Kerr effect. However, currently the non-Gaussian effects induced by known non-linear optical materials are extremely weak RevModPhys.77.513 ; Slussarenko2019 ; Kok2007 ; Ladd2010 ; PhysRevA.100.052301 , and the generation of arbitrary non-Gaussian operations is difficult. To circumvent these difficulties an alternative is the measurement-based approach (see References PhysRevA.100.012326 ; PhysRevA.100.052301 and references therein), which we will be using here. Here non-Gaussianity is introduced by a process of post-selected measurement on entangled qumodes (in other words the accepted photons are filtered by the result of the measurement on an entangled ancilla qumode).
This Section will present the necessary gate operations of both kinds that will be required for our study. In Section 2.1, we review the action of the various Gaussian gate operations, and then Section 2.2 outlines the construction of non-Gaussian operators. Although the methods presented here are general, this study will utilise the Gaussian gate operations available on the StrawberryFields platform Killoran_2019 ; Bromley_2020 .
@@ 2.1 Review of required Gaussian gate operations
The required Gaussian gate operations in the continuous-variable regime are expressed either in terms of the quadrature operators, $\hat{x}$ and $\hat{p}$, or the creation and annihilation operators, $\hat{a}^{\dagger}$ and $\hat{a}$, of the conventional SHO defined in Equation 3. Here we will outline the gates by highlighting the action that is most relevant for this study (for a full set of expressions the reader is referred to Reference RevModPhys.77.513 ):
@@
    As the phase $e^{-i\theta/2}$ corresponding to the ground state energy acts universally, it will usually be possible to ignore it.
  • Controlled-X: We will in addition to the above single qumode operators be using several Gaussian two-qumode operators. These induce displacements in $x$ or $p$ of qumode-$x$ which depend on the value of $y$ or $p_{y}$ measured on a second qumode-$y$ to which it is coupled, and vice-versa. In terms of operators the controlled-X gate (again taking $\hbar=1$) for two qumodes with variables $x,p_{x},y,p_{y}$ is
@@
[…]
As a warm-up exercise, let us perform a few test manipulations utilising some of these gates on a qumode groundstate. Figure 0(a) shows the groundstate $\left\langle x|0\right\rangle$ together with two manipulations corresponding to $S(\log(1/2))\left\langle x|0\right\rangle$ which produces a widened wave function (in yellow), and a second state corresponding to
$S(\log(4))S(\log(1/2))\left\langle x|0\right\rangle$
@@
and then in this example we perform a homodyne measurement at $y=1$ resulting in the displaced state $\left\langle x-s|0\right\rangle$. The first case (yellow line) takes $s=2$ resulting in $\psi(x)\to\psi(x-2)$ and the curve moves to the right. The second case (orange line) repeats the displacement operation with $s=-4$ resulting in the displacement $\psi(x-2)\to\psi(x+2)$ and the curve is then displaced to the left by 4 units. The two sets of manipulations which produced Figures 0(a) and 0(b) were generated respectively by the two simple circuit diagrams shown in Figures 1(a) and 1(b).

### v2

@@ Abstract
We delve into the use of photonic quantum computing to simulate quantum mechanics and extend its application towards quantum field theory. We develop and prove a method that leverages this form of Continuous-Variable Quantum Computing (CVQC) to reproduce the time evolution of quantum-mechanical states under arbitrary Hamiltonians, and we demonstrate the method’s remarkable efficacy with various potentials. Our method centres on constructing an evolver-state, a specially prepared quantum state that induces the desired time-evolution on the target state. This is achieved by introducing a non-Gaussian operation using a measurement-based quantum computing approach, enhanced by machine learning. Furthermore, we propose a framework in which these methods can be extended to encode field theories in CVQC without discretising the field values, thus preserving the continuous nature of the fields. This opens new avenues for quantum computing applications in quantum field theory.
@@ 1 Introduction
Harnessing the intricate dynamics of quantum mechanics to improve our understanding of fundamental physics has led to the pursuit of computational paradigms that go beyond classical boundaries and allow for complex systems to be simulated with devices which themselves are inherently quantum mechanical. Such devices would leverage the properties of quantum systems, such as superposition and entanglement, to perform calculations directly using the intrinsic dynamics of the system. Two main paradigms have been identified: digital and analogue quantum computing. The former encodes information onto systems with a finite number of discrete degrees of freedom, such as the qubit, a two-state quantum system. The latter instead encodes information on systems which are described by operators which have a continuous spectrum. This paradigm is often called Continuous-Variable Quantum Computing (CVQC) Lloyd1999 ; RevModPhys.77.513 ; Ladd2010 ; Adesso_2014 . Among the various kinds of CVQC, quantum optics emerges as a fascinating framework, using the infinite-dimensional landscape of photon states to encode and manipulate information PhysRevA.73.032318 ; Kok2007 ; Slussarenko2019 ; Killoran_2019 ; Bromley_2020 ; Eaton_2022 ; Taballione2023modeuniversal . The eigenstates of these operators form an infinite-dimensional Hilbert space, with the continuous-variable analog of the qubit being the so-called qumode.
@@
CVQC, rooted in the continuous spectra of quantum operators, offers various possibilities for simulating the dynamics of quantum particles and fields. By employing such qumodes — quantum analogues of classical harmonic oscillators — as the fundamental information units, CVQC provides a natural framework for encoding quantum states in the continuous observables of photons, e.g. their position or momentum. This paradigm allows for implementing Gaussian gate operations, which manipulate the quantum states through transformations that preserve their Gaussian character, thereby enabling a broad range of quantum simulations. However, the true power of CVQC unfolds with the inclusion of non-Gaussian operations, which introduce higher-order interactions essential for achieving universal quantum computation Lloyd1999 ; PhysRevA.100.012326 ; PhysRevA.100.052301 . These operations, albeit challenging to implement due to the weakly interacting nature of photons, open the door to simulating complex quantum systems with high fidelity. Most existing proposals to simulate quantum systems on CVQC rely on specific ways to induce non-Gaussian effects, such as the Kerr effect. However, the non-Gaussian characteristics introduced by current non-linear optical materials are very weak RevModPhys.77.513 ; Slussarenko2019 ; Kok2007 ; Ladd2010 ; PhysRevA.100.052301 , and constructing an arbitrary non-Gaussian operation is difficult. An alternative approach is achieved by integrating measurement-based quantum computing techniques PhysRevLett.86.5188 ; Booth:2021hvw and leveraging the entanglement of qumodes. Through this, it is possible to instead induce the desired non-Gaussian characteristics, paving the way for simulations that capture non-trivial quantum dynamics.
@@
A central aspect of quantum systems that one might wish to explore using such methods is the Hamiltonian and the time-evolution that is governed by it, a cornerstone in understanding the dynamics of quantum particles and fields Jordan:2011ci ; Jordan:2012xnu ; Jordan:2014tma ; PhysRevA.92.063825 ; Jordan_2018 ; Klco:2018zqz ; Banuls:2019bmf . By simulating the time evolution governed by a system’s Hamiltonian, we can explore how quantum states change over time, which is crucial for predicting the behaviour of quantum particles and systems under various conditions. This process is essential for simulating inherently quantum-mechanical phenomena that cannot be accurately modelled using classical physics. One of the most notable examples is quantum tunnelling, a phenomenon where particles pass through potential barriers that would be insurmountable, according to classical mechanics. This process is critical in a wide range of quantum systems, from the decay of atomic nuclei to the operation of quantum dots and superconducting qubits. By simulating the time evolution of quantum systems, one can also investigate other observables, such as energy spectra, correlation functions, and phase transitions, providing deep insights into the nature of quantum materials, chemical reactions, and even the evolution of early universe conditions in high-energy physics. CVQC offers a novel approach to simulating the time evolution of quantum states under arbitrary Hamiltonians. By decomposing the time evolution into discrete steps through Trotterization, we show how CVQC facilitates the simulation of complex quantum systems, including those governed by non-Gaussian potentials.
@@
In contrast, quantum gate computing operates within a digital framework, encoding quantum information in discrete qubits and manipulating it through a sequence of quantum gates. While this approach has paved the way for significant advancements in quantum computing, and has gained interest for uses in the simulation of quantum field theories Jordan:2011ci ; Jordan:2012xnu ; Jordan:2014tma ; PhysRevA.92.063825 ; Jordan_2018 ; Klco:2018zqz ; Banuls:2019bmf ; Abel:2020qzm ; Abel:2020ebj ; PRXQuantum.4.027001 ; Kane:2022ejm , high-energy particle collisions Blance:2020ktp ; Bepari:2020xqi ; Bepari:2021kwv ; Gustafson:2022dsq , and machine learning Schuld2015 ; Carleo:2019ptp ; Abel:2022lqr ; Abel:2023erv ; Rousselot:2023pcj , it inherently approximates the continuous nature of quantum systems, potentially limiting its ability to capture the full spectrum of quantum dynamics. Thus, CVQC, with its continuous-variable approach, offers a promising avenue for simulating quantum systems complementary to quantum gate computing approaches.
@@
The simulation of non-trivial Hamiltonians on CVQC devices has been difficult due to the challenges in implementing non-Gaussian operations on photonic devices. Most current approaches propose circuits which involve non-Gaussian gate operations generated by non-linear optics to achieve Hamiltonian simulation on a continuous-variable device PhysRevA.92.063825 ; PhysRevA.97.062311 . However, experimentally such operations are difficult to produce Lloyd1999 ; RevModPhys.77.513 ; Adesso_2014 . This has led to attempts to generate non-Gaussian effects through the use of Gaussian operations and measurements PhysRevA.105.012412 , in particular References PhysRevA.100.012326 ; PhysRevA.100.052301 utilise a machine-learning routine to enhance the production of non-Gaussian states. In this paper, we build and improve on this approach and propose a quantum circuit for simulating the Trotterised time-evolution of a quantum-mechanical state under the influence of an arbitrary Hamiltonian using only Gaussian operations and measurements. The circuit for a single Trotter step can then be applied iteratively to achieve the time-evolution of the Hamiltonian. By using a top-hat resource function and generating the evolver-state, the non-Gaussian part of the Trotter evolution, using a measurement-based circuit, we show that the effect of the noise-factor from Reference PhysRevA.100.012326 can be maximally suppressed without reducing the strength of the overall operation. We demonstrate the circuit’s ability to simulate time-evolution for several examples of quantum-mechanical Hamiltonians and find that the circuit performs remarkably well compared to exact, classical simulations. Furthermore, we show that the approach can be extended, outlining how the continuous behaviour of the CVQC device can be harnessed to accurately simulate quantum field theories, maintaining the continuous nature of fields.
@@ 2 Background: Continuous-Variable Quantum Computing
The framework we will consider is quantum computation via quantum optics, which offers an experimentally realisable approach to Continuous-Variable Quantum Computing (CVQC), in which the continuous-variable system is constructed from the quantised electromagnetic field RevModPhys.77.513 ; Adesso_2014 . The use of quantum optics as a method of quantum computing benefits from the exceedingly low decoherence properties of photons, and therefore holds good potential for transmitting and maintaining quantum information throughout many gate operations Kok2007 ; Ladd2010 ; Bromley_2020 . In this realisation, the Hilbert space of each qumode is the infinite-dimensional photon-number degree of freedom EISERT2003 which is the Fock basis, such that the total Hilbert space, $\mathscr{H}$, of an $N$ qumode system is
@@
$[\hat{x}_{i},\hat{p}_{j}]=i\hbar\delta_{ij}.$ (4)
It will be convenient to express the state on a qumode as an expansion in the Fock basis, such that
$|\psi\rangle~{}=~{}\sum^{\infty}_{n=0}A_{n}|n\rangle~{},$ (5)
where $A_{n}$ is the coefficient of the $n$ th Fock state, $\left|n\right\rangle$. For the rest of this paper, we will adopt natural units, $\hbar=m=\omega=1$.
@@
where $i$ and $j$ label the qumodes on which they act, and $\theta_{ij},\theta^{\prime}_{ij}$ and $\theta^{\prime\prime}_{ij}$ are constants. For universal computation it is essential also to be able to implement non-Gaussian gates (i.e. with $\hat{x}^{3}$ and higher appearing in the phase) Lloyd1999 ; PhysRevA.100.012326 ; PhysRevA.100.052301 . This is a delicate process because the fact that photons do not strongly interact with each other, while being excellent for maintaining coherence, is accompanied by the negative implication that non-Gaussian effects are hard to achieve. Directly producing non-Gaussian gate operations requires the use of non-linear optical materials which induce non-Gaussian effects, such as the Kerr effect. However, currently the non-Gaussian effects induced by known non-linear optical materials are extremely weak RevModPhys.77.513 ; Slussarenko2019 ; Kok2007 ; Ladd2010 ; PhysRevA.100.052301 , and the generation of arbitrary non-Gaussian operations is difficult. To circumvent these difficulties an alternative is the measurement-based approach (see References PhysRevA.100.012326 ; PhysRevA.100.052301 and references therein), which we will be using here. Here non-Gaussianity is introduced by a process of post-selected measurement on entangled qumodes (in other words the accepted photons are filtered by the result of the measurement on an entangled ancilla qumode).
This Section will present the necessary gate operations of both kinds that will be required for our paper. In Section 2.1, we review the action of the various Gaussian gate operations, and then Section 2.2 outlines the construction of non-Gaussian operators. Although the methods presented here are general, this paper will utilise the Gaussian gate operations available on the StrawberryFields platform Killoran_2019 ; Bromley_2020 .
@@ 2.1 Review of required Gaussian gate operations
The required Gaussian gate operations in the continuous-variable regime are expressed either in terms of the quadrature operators, $\hat{x}$ and $\hat{p}$, or the creation and annihilation operators, $\hat{a}^{\dagger}$ and $\hat{a}$, of the conventional SHO defined in Equation 3. Here we will outline the gates by highlighting the action that is most relevant for this paper (for a full set of expressions the reader is referred to Reference RevModPhys.77.513 ):
@@
    As the phase $e^{-i\theta/2}$ corresponding to the ground-state energy acts universally, it will usually be possible to ignore it.
  • Controlled-X: We will in addition to the above single qumode operators be using several Gaussian two-qumode operators. These induce displacements in $x$ or $p$ of qumode-$x$ which depend on the value of $y$ or $p_{y}$ measured on a second qumode-$y$ to which it is coupled, and vice-versa. In terms of operators the controlled-X gate (again taking $\hbar=1$) for two qumodes with variables $x,p_{x},y$ and $p_{y}$ is
@@
Figure 1: Example operations using a squeezing gate, and a controlled-X gate together with homodyne measurement. In (a) the ground state is squeezed by $S(\ln(1/2))$ which produces a wave-function flattened by a factor of 2 (yellow line). Then it is squeezed again by $S(\ln\,(4))$ producing a ground state squeezed by a factor 2 (orange line). In (b) we perform a composite displacement, by using a Controlled-X gate followed by a homodyne measurement of $y$, first displacing by 2 to the right, so that $\psi(x)\to\psi(x-2)$ (yellow line), then displacing by 4 to the left, so that $\psi(x-2)\to\psi(x+4)$ (red line). In this case the Fock truncation is 60, and we see some distortion beginning to appear at the peak.
As a warm-up exercise, let us perform a few test manipulations utilising some of these gates on a qumode groundstate. Figure 1(a) shows the groundstate $\left\langle x|0\right\rangle$ together with two manipulations corresponding to $S(\ln(1/2))\left\langle x|0\right\rangle$ which produces a widened wave function (in yellow), and a second state corresponding to
$S(\ln(4))S(\ln(1/2))\left\langle x|0\right\rangle$
@@
and then in this example we perform a homodyne measurement at $y=1$ resulting in the displaced state $\left\langle x-s|0\right\rangle$. The first case (yellow line) takes $s=2$ resulting in $\psi(x)\to\psi(x-2)$ and the curve moves to the right. The second case (orange line) repeats the displacement operation with $s=-4$ resulting in the displacement $\psi(x-2)\to\psi(x+2)$ and the curve is then displaced to the left by 4 units. The two sets of manipulations which produced Figures 1(a) and 1(b) were generated respectively by the two simple circuit diagrams shown in Figures 2(a) and 2(b).